Prove that cos6x = 32cos^6x - 48cos^4x + 18cos^2x – 1

Grade:11

1 Answers

Arun
6 years ago
cos6x=cos[2(3x)]=2cos2(3x)-1 cos2x=2cos2x-1 =2[cos(3x)]2-1 cos3x=4cos3x-3cosx =2[4cos3x-3cosx]2-1 (a-b)2=a2+b2-2ab =2[16cos6x+9cos2x-24cos4x]-1 =32cos6x-48cos4x+18cos2x-1